Dinitrogen oxide, N2O, which is used in dentistry, is an anesthetic also called laughing gas. What pressure, in atmospheres, of 0.350 mole of N2O at the 22°C in a 5.00-L container?


1
Expert's answer
2021-11-15T10:50:45-0500

Pressure of N2O = nRT/V

= (0.35×0.0821×295)/5 = 1.7 atm